# LLMs in the SOC: An Empirical Study of Human-AI Collaboration in Security Operations Centres

### v1

@@ 1 Introduction
Recently, LLMs have emerged as promising tools to support such SOC workflows, with applications ranging from threat intelligence to triage automation [7, 8, 9, 10]. While prior research explores human-LLM interaction in various domains and lab studies [11, 12, 13, 14], it does not capture how SOC analysts utilise LLMs in their daily workflows. Existing work, such as Copilot Guided Response (CGR) [7], show system-level performance, not individual analysts’ usage patterns.
Gaining insights into how SOC analysts utilise LLMs in practice, including the types of questions they ask and the tasks they use them for, is essential for evaluating the practical role of LLMs in operational cybersecurity (Figure 1). This study addresses that gap by analysing 3,090 queries submitted to an internally deployed instance of GPT-4 over 10 months (May 2023 to March 2024) by 45 SOC analysts at eSentire Inc., a cybersecurity firm specialising in 24/7 managed detection and response. To guide this investigation, we focus on the following research question:
@@
The answers to this question reveal SOC analysts’ task priorities and engagement patterns with LLMs, informing the design of more effective LLM-based assistants as human-AI collaboration becomes increasingly common in SOC environments. We address this question through a multi-level analysis of analyst-LLM interactions across queries, tasks, and conversations.
@@ 2 Related Works
2.1 LLMs for Cybersecurity
@@
Prior to LLMs, SOCs already leveraged AI across SIEM correlation, IDS/IPS, and related tooling [17]. Recent works highlights the expanding role of LLMs in cybersecurity workflows. For example, Copilot Guided Response (CGR) aids triage and remediation by generating recommendations from historical threats [7], while tools like LocalIntel [18] extract structured threat intelligence from unstructured sources to accelerate threat contextualisation. LLMs also support vulnerability scanning [19, 20, 21], anomaly detection [22, 23], intrusion detection[24], phishing simulation [25, 26, 27], threat intelligence summarisation [28] and privacy-preserving analysis [29], leveraging their generalisation and pattern recognition capabilities. However, most of these systems are evaluated in controlled environments, with a focus on performance metrics such as speed and accuracy of decision-making (e.g. [7]). In contrast, our study provides the first empirical analysis of LLM usage in a live SOC, examining how analysts actually engage with the LLMs, what help they seek, how interactions unfold, and how adoption evolves over time, offering a unique, human-centred view into analyst-LLM interactions in practice.
2.2 LLM Roles in Decision Support
@@
LLMs play diverse roles in human-AI decision-making, shaped by system autonomy, user needs, and interaction structure [30, 31, 32, 33, 34]. LLMs span roles from assisted intelligence (retrieval and summarisation) to more collaborative forms like augmented and cooperative intelligence, which aid reasoning and joint problem structuring [30, 33]. HCI research highlights the importance of aligning AI support with decision phases and initiatives, whether recommending actions, making predictions, proposing alternatives, or surfacing evidence [35, 32, 31, 33]. While AI suggestions can speed decisions, they may also increase over-reliance [36, 37]. Cognitive forcing techniques (e.g., delaying AI output) can reduce over-reliance but risk under-reliance among experienced users [36, 38]. Evaluative AI offers a middle ground by surfacing pros and cons without prescribing decisions [39], supporting autonomy in machine-in-the-loop systems [40, 41, 42]. Our analysis examines how LLMs operate in practice and the types of decision support that analysts seek.
2.3 Interaction Patterns and Analyst Strategies
@@
Despite the growing capabilities of models, interaction patterns between humans and LLMs remain underexplored. Studies show that interaction framing and modality shape user expectations, trust, and behaviour [13, 43, 12, 11, 44, 45, 27]. Recent work highlights that prompt structure shapes explainability and collaboration [13], dialogue in complex tasks involves verbal and nonverbal cues [44], and human–AI interaction spans planning to testing across varying agent roles [11]. These studies underscore that the effective integration of LLMs for decision-making hinges on their ability to flexibly assume different roles shaped by user needs, interaction design, and evolving task context. Our study addresses this gap by analysing thousands of analyst-LLM interactions in an operational SOC, uncovering concrete interaction patterns, evolving strategies, and the situated use of LLMs as cognitive aids in high-stakes, time-sensitive environments.
@@ 3 Method
To analyse and understand SOC analysts’ interactions with LLMs, we adopted a six-phase methodology, as illustrated in Figure 2, which we describe in detail below.
@@
To ensure consistency and reliability in applying this coding scheme, a sample of 20 question-LLM response pairs was independently coded by all three researchers. This was followed by a calibration session where the same 20 pairs were jointly reviewed and coded to to reconcile differences and refine definitions. Once alignment was achieved, the researchers independently coded a further 310 (10%) question-response pairs to test their collective understanding of the attributes and coding scheme. The reconciliation sessions helped resolve any disagreements and update the codebook. IRR, measured using Fleiss’ Kappa across the three dimensions, showed strong agreement: 0.90 for Query Pattern, 0.82 for Query Subject, and 0.79 for Task, indicating substantial to near-perfect agreement.
@@
Conversation tagging. The next step involved grouping successive analyst queries that relate to the same topic or investigative thread and treating them as a single conversational unit by considering (a) overlapping security artifacts (e.g., commands, email rules), (b) continuity in investigative intent (e.g., both steps probing the same log entry), and (c) occurring within a 30-minute window (used only as a guideline rather than a strict cutoff). Each conversation was assigned a unique identifier (e.g., C1, C2). To ensure consistency and reliability in tagging, the three researchers independently labelled a random sample of 45 analyst questions, comprising 142 question–response pairs, the met to reconcile and refine their understanding of the tagging process. Once consistent tagging practices were established, an additional 240 instances (13%) were tagged as a further verification step. Fleiss’ Kappa was used to calculate IRR, resulting in a score of 0.75, indicating substantial agreement.
@@
Figure 2: The six-phased approach to analyse and understand SOC analysts’ interactions with LLMs.
@@
Phase 3: Remaining Dataset Question Coding & Conversation Tagging. In this phase, given acceptable agreement in Phase 2, the remaining question-LLM response pairs were divided among the three researchers, with each coding approximately 900 pairs independently. After initial coding, each researcher’s annotations were cross-verified by another team member to ensure consistency and accuracy; discrepancies were resolved through discussions. To further enhance the reliability of the coding process, one researcher conducted a comprehensive review of all the coded data. Discrepancies were minimal, observed only in 4% of the cases and resolved collaboratively through discussions.
@@
Phase 4: Query Level Thematic Analysis. One researcher consolidated all analyst question-LLM response pairs into a dataset and conducted query level thematic analysis [15], as outlined in Figure 2. This phase built on earlier manual coding (Phase 3) and introduced a hybrid approach that combined qualitative coding with data-driven clustering and thematic interpretation (Phase 5) to surface higher-level insights. We applied semantic clustering to the coded data to support theme development. Sentence-BERT embeddings (stsb-roberta-large) and agglomerative clustering with cosine similarity [46] were used to group over 1000 distinct query patterns into semantically similar codes, supporting pattern abstraction. All clusters were manually reviewed to ensure semantic coherence. Similar patterns such as explain [x] and explain this [x] were grouped. Themes were then developed by merging related clusters and assigning meaningful, high-level labels spanning across all attributes; e.g., command and multiple commands formed a broader command-related subject theme.
@@
To ensure the robustness of these themes, we conducted a triangulation process [47] in which two additional researchers, who were not involved in the clustering, independently reviewed and refined the themes. This hybrid approach, combining manual coding, computational clustering, and human interpretation, ensured both methodological rigour and interpretive depth [48, 49].
@@ 4 Results
Analyst and Model Information
The 45 SOC analysts comprised: 1) SOC Analyst I: 52%; 2) Senior SOC Analyst: 17%; 3) Threat Analyst: 14%; 4) SOC Analyst II: 8%; 5) SOC Incident Handler: 2%; 6) Associate Analysts: 7%. All participants, except Associate Analysts, had 3–5 years of experience and worked with a proprietary SOC stack integrating alerts, logs, and threat intelligence. They used GPT-4 (GPT-4-0613) was used during live investigations without fine-tuning or access to the internet or SOC systems. Although other models were available, our analysis focuses solely on GPT-4. Analysts were free to use the LLM as they saw fit, but were instructed not to share sensitive or identifying data.
Figure 3 (left) summarises analyst engagement with the LLM over 10 months. Usage was uneven, with a small number of analysts generating the majority of queries. Given the skewed usage, we clustered the analysts based on their usage before analysing their long-term reliance on the LLM. We applied the Fisher-Jenks natural breaks algorithm [^1] https://en.wikipedia.org/wiki/Jenks_natural_breaks_optimization to total query counts per analyst. This generated three distinct clusters: a large group of low-usage analysts (Cluster 0, $n=34$), a smaller group of moderate users (Cluster 1, $n=10$), and a single power user (Cluster 2, $n=1$; most active analyst). Figure 3 (right) shows that daily usage increased steadily over time, driven by a subset of highly active analysts, suggesting deeper integration of the LLM into daily workflows. The number of active analysts also grew month on month, suggesting sustained interest and continuous adoption of the LLM.
@@ 4.1 Exploratory Analysis
While a single power user submitted approximately 600 queries (17% of the total dataset), usage was not limited to a few individuals. By early 2024, over 45 analysts had interacted with the system, and daily query volumes increased from fewer than 10 queries per day in May-August 2023 to over 30 queries per day in January and February 2024. We analysed the monthly usage patterns and observed that the number of active analysts grew from 4 in May 2023 to 13 in June, reaching 18 by July and steadily increasing to a peak of 25 in February 2024. In just the first few days of March 2024, 16 analysts had already submitted queries, indicating continued high engagement.
@@ 4.2 Query Level Analysis
This section discusses the thematic analysis results, which involved coding the individual queries. Note that each conversation could involve multi-step interactions. The thematic analysis was conducted on individual queries, rather than entire conversations, to avoid missing the granular insights these provide over a holistic understanding of conversations. Section 4.3 presents conversation-level analysis. A complete list of themes can be found in the appendix.
The findings from the exploratory analysis reveal an interesting landscape of LLM use in our dataset. Despite wide variation in query volume, engagement style, and query lengths, usage trends point to a transition from periodic exploration to routine integration, with increased adoption over time and sustained interaction by a subset of analysts stronger increase among 11/45. These observations offer grounded insights into how LLMs are appropriated in practice, setting the stage for the next phase of our study: a qualitative analysis of the tasks and interaction strategies underpinning analyst-LLM collaboration.
@@ 4.2.1 SOC Analyst Task Types
Task themes reflect the actual work that analysts aim to accomplish when interacting with an LLM. We find that analysts used LLMs for a range of tasks central to day-to-day SOC operations [1, 3]. These include interpreting and analysing commands; editing and improving the clarity of incident descriptions, documentation, and alert text; and understanding or troubleshooting code, scripts, and regular expressions. Other common uses included analysing files and malware, exploring tool and system behaviours, constructing queries for data lookup, and clarifying email rules or suspicious artifacts, highlighting the breadth of tasks analysts juggle in operational contexts. Below, we detail three of the most common task themes (see Figure 4 for the top 10) illustrating how LLMs are integrated in operational workflows (see Figure 9 in the appendix to observe how task adoption changed over time).
@@
The LLM often began with an overview of the command, explaining the components and their relationships, inferring the intent behind complex operations, and highlighting suspicious behaviours. For example:
“In summary, the [anonymised command] is trying to login as [anonymised user] without password and $\ldots$
Many responses also included parameter breakdowns, example outcomes, risk implications, and recommended follow-up actions. Based on their sustained usage, we hypothesise that the LLM helped by decomposing command components and synthesising likely intent and behaviour, , the LLM has shown the potential of enhancing the analysts’ situational understanding. This aligns with prior work (e.g., [16]), which finds that SOC practitioners prefer explanations that contextualise evidence and inform action.
@@
In addition to the top three themes, analysts used LLMs for a wide range of other tasks including Tool/ Software/ System Understanding; Attack, Security & Threat Intelligence; Access Control & Authentication; File & Malware Analysis; Network & Traffic Analysis; and Query, Search & Data Processing. These tasks reflect specialised needs where LLMs provide targeted support. Notably, LLM Chat interactions suggest a growing social relationship between analysts and the assistant; five analysts explicitly thanked the LLM.

### v2

@@ 1 Introduction
Recently, LLMs have emerged as promising tools to support such SOC workflows, with applications ranging from threat intelligence to triage automation [7, 8, 9, 10]. While prior research explores human-LLM interaction in various domains and lab studies [11, 12, 13, 14], it does not capture how SOC analysts utilise LLMs in their daily workflows. Existing work, such as Copilot Guided Response (CGR) [7], shows system-level performance, not individual analysts’ usage patterns.
Gaining insights into how SOC analysts utilise LLMs in practice, including the types of questions they ask and the tasks they use them for, is essential for evaluating the practical role of LLMs in operational cybersecurity (Figure 1). This study addresses that gap by analysing 3,090 queries submitted to GPT-4 during live investigations over 10 months (May 2023 to March 2024) by 45 SOC analysts at eSentire Inc., a cybersecurity firm specialising in 24/7 managed detection and response [15]. To guide this investigation, we focus on the following research question:
@@
The answers to this question reveal the task priorities and engagement patterns of SOC analysts with LLMs, informing the design of more effective LLM-based assistants as human-AI collaboration becomes increasingly common in SOC environments. We address this question through a multi-level analysis of analyst-LLM interactions across queries, tasks, and conversations.
@@ 2 Related Works
AI for SOCs. SOCs integrate multiple systems including SIEM platforms that ingest telemetry from IDS/IPS, firewalls, and EDR tools; SOAR platforms for response coordination; and supplementary tools such as User and Entity Behaviour Analytics (UEBA) and Threat Intelligence Platforms (TIPs) [18, 19]. These systems have long relied on AI to enhance operational efficiency and decision-making. Prior to the emergence of LLMs, AI was already embedded across the SOC workflow, from threat detection [20], alert triage [21], and incident response [22] to end-to-end automation [19]. SIEMs employed ML/DL for event correlation, anomaly detection, and prioritisation [18], while TIPs used NLP and graph inference to enrich contextual understanding [19]. Commercial tools such as Microsoft Sentinel, Maltego, and Splunk exemplify this integration, offering AI-driven capabilities for visualisation, triage, and orchestration [7, 18, 19, 23, 6]. While effective, these approaches were task-specific and limited in their ability to reason across diverse, unstructured contexts. The recent emergence of LLMs provides a broader, context-aware form of cybersecurity reasoning, enabling more adaptive and human-aligned support across SOC operations [23].
@@
LLMs for Cybersecurity. Broadly, LLMs have been used to support vulnerability scanning, anomaly/intrusion detection, phishing simulation, threat intel summarisation, and privacy-preserving analysis [24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. Recent works highlight the expanding role of LLMs in cybersecurity workflows to support SOC analysts. For example, Copilot Guided Response (CGR) aids triage and remediation by generating recommendations from historical threats [7]. Most of these systems are evaluated in controlled environments, with a focus on performance metrics such as speed and accuracy of decision-making. In contrast, our study offers the first empirical analysis of LLM use in a live SOC, examining how analysts engage with LLMs, what help they seek, how interactions unfold, and how adoption evolves, providing a unique human-centred view of analyst-LLM practice.
@@
LLM Roles in Decision Support. LLMs play diverse roles in human-AI decision-making, shaped by system autonomy, user needs, and interaction structure [34, 35, 36, 37, 38]. LLMs span roles from assisted intelligence (retrieval and summarisation) to more collaborative forms like augmented and cooperative intelligence, which aid reasoning and joint problem solving [34, 37]. HCI research stresses aligning AI support with decision phases, whether recommending actions, predicting, proposing alternatives, or surfacing evidence [39, 36, 35, 37]. While AI suggestions can speed decisions, they may also increase over-reliance [40, 41]. Cognitive forcing techniques (e.g., delaying AI output) can reduce over-reliance but risk under-reliance [40, 42]. Evaluative AI provides balanced insights, highlighting pros and cons without dictating choices, thus supporting autonomy in machine-in-the-loop systems [43, 44, 45, 46]. We examine how LLMs function in practice and what decision support analysts seek.
Interaction Patterns and Analyst Strategies.
@@
Despite the growing capabilities of models, interaction patterns between humans and LLMs remain underexplored. Studies show that interaction framing and modality shape user expectations, trust, and behaviour [13, 47, 12, 11, 48, 49, 31]. Recent work highlights that prompt structure shapes explainability and collaboration [13], dialogue in complex tasks involves verbal and nonverbal cues [48], and human–AI interaction spans planning to testing across varying agent roles [11]. These studies show that effective LLM integration depends on their ability to adapt to user needs, interaction design, and task context. Our study addresses this gap by analysing thousands of analyst-LLM interactions in an operational SOC, uncovering concrete interaction patterns, evolving strategies, and the situated use of LLMs as cognitive aids in high-stakes, time-sensitive environments.
@@ 3 Method
To analyse and understand SOC analysts’ interactions with LLMs, we adopted a five-phase methodology, as illustrated in Figure 2, which we describe in detail below. This approach is guided by Nowell et al.’s [16] multi-phase framework for thematic analysis, encompassing: (i) familiarisation with data (phases 1-2), (ii) code generation (phases 2-3), (iii) theme identification, review, and definition (phase 4), and (iv) reporting (phase 5). We follow their iterative process and incorporate specific trustworthiness practices, including multi-coder independent coding with strong inter-rater reliability (IRR), triangulation, reconciliation, cross-verification, and systematic documentation.
@@
To ensure consistent application of the agreed coding scheme, all three researchers independently coded a sample of 20 question-LLM response pairs. This sample was used to verify their initial shared understanding of coding the three attributes. This was followed by a calibration session, during which the same 20 pairs were jointly reviewed and coded to reconcile differences, refine the definitions, and develop a preliminary codebook. Once alignment was achieved, the researchers independently coded a further 310 (10%) question-response pairs to further test their collective understanding of the attributes and coding scheme before being confident in independently coding the remaining pairs. IRR, measured using Fleiss’ Kappa across the three dimensions, showed strong agreement: 0.90 for Query Pattern, 0.82 for Query Subject, and 0.79 for Task, indicating substantial to near-perfect agreement. A reconciliation session helped resolve any disagreements and update the codebook.
@@
Conversation tagging. Following a similar process described above, the next step involved grouping successive analyst queries that relate to the same topic or investigative thread and treating them as a single conversational unit by considering (a) overlapping security artifacts (e.g., commands, email rules), (b) continuity in investigative intent (e.g., both steps probing the same log entry), and (c) occurring within a 30-minute window (used only as a guideline rather than a strict cutoff). Each conversation was assigned a unique identifier (e.g., C1, C2). To ensure consistency and reliability in tagging, the three researchers independently labelled a random sample of 45 analyst questions, comprising 142 question–response pairs, then met to reconcile and refine their understanding of the tagging process. Once consistent tagging practices were established, an additional 240 instances (13%) were tagged as a further verification step. Fleiss’ Kappa was used to calculate IRR, resulting in a score of 0.75, indicating substantial agreement. Differences were reconciled through discussions.
@@
Figure 2: The five-phased approach to analyse and understand SOC analysts’ interactions with LLMs.
@@
Phase 3: Remaining Dataset Question Coding & Conversation Tagging. In this phase, given acceptable agreement in Phase 2, the remaining question-LLM response pairs were divided among the three researchers, with each coding approximately 900 pairs independently. After initial coding, each researcher’s annotations were cross-verified by another team member to ensure consistency and accuracy; discrepancies were resolved through discussions. To further enhance the reliability of the coding process, the lead author conducted a comprehensive review of all the coded data. Discrepancies were minimal, observed only in 4% of the cases and resolved collaboratively through discussions.
@@
Phase 4: Query Level Thematic Analysis. One researcher consolidated all analyst question-LLM response pairs into a dataset and conducted query-level thematic analysis [50], as outlined in Figure 2. This phase built on earlier manual coding (Phase 3) and introduced a hybrid approach that combined qualitative coding with data-driven clustering and thematic interpretation (Phase 5) to surface higher-level insights. We applied semantic clustering to the coded data to support theme development. Sentence-BERT embeddings (stsb-roberta-large) and agglomerative clustering with cosine similarity [51] were used to group over 1000 distinct query patterns into semantically similar codes, supporting pattern abstraction. All clusters were manually reviewed to ensure semantic coherence, considering the original query and LLM response. Similar patterns such as explain [x] and explain this [x] were grouped. Themes were then developed by merging related clusters and assigning meaningful, high-level labels spanning across all attributes; e.g., command and multiple commands formed a broader command-related subject theme. For example, semantic clustering was applied to cluster 1,354 initial query pattern codes into 980 unnamed clusters, which were iteratively merged by the lead author into 147 codes and finally grouped into themes.
@@
To ensure the robustness of these themes, we conducted a triangulation process [16] in which two researchers, who were not involved in the clustering, independently reviewed and refined the themes. This hybrid approach, combining coding, clustering, and human insight, balanced rigour with interpretive depth. [52, 53].
@@ 4 Results
Analyst Role Count Responsibilities
SOC Analyst I 24 Handles initial alert triage and prioritisation, basic malware and network traffic analysis, IP/domain reputation checks, OSINT research, incident documentation, false positive tuning, and escalates complex cases.
SOC Analyst II 4 Coordinates intermediate incident response, manages specialised security tools, performs advanced log analysis and digital forensics, develops automation scripts, liaises with vendors, and reviews junior analysts’ work.
Senior SOC Analyst 7 Leads advanced investigations and root cause analysis, conducts malware reverse engineering, runs threat hunting campaigns, mentors staff, optimises SIEM rules, coordinates across teams, and drives process improvements.
SOC Incident Handler 1 Collects and analyses threat intelligence, develops IOCs, attributes threat actors, monitors threat landscape, produces reports for leadership, integrates external feeds, and performs strategic risk assessments.
Threat Analyst 6 Oversees major incident response, containment, and eradication, handles stakeholder communication and coordination, facilitates post-incident reviews, executes emergency procedures, and supports continuity planning.
Associate Analyst 3 Provides entry-level monitoring, basic research, alert categorisation, report generation, and administrative support, while engaging in training and certification programs.
TABLE I: Roles and responsibilities of the 45 participating SOC analysts.
@@
4.1 Participant Overview and LLM Usage Constraints
Participants. Table I summarises the roles and responsibilities of the 45 SOC analysts who participated in the study. The majority were Tier 1 (SOC Analyst I), reflecting their front-line role in alert triage. Tier 2 (SOC Analyst II) and Tier 3 (Threat Analyst, Incident Handler) roles are more specialised and fewer in number, accounting for their lower representation. All participants, except for Associate Analysts (interns/co-ops, n=3), had 3–5 years of experience. Participation in the study was voluntary and the distribution of Tier 1, Tier 2, and Tier 3 analysts is not representative of actual staffing ratios. Analysts opted into the LLM pilot study via an internal Slack announcement, which invited exploratory use of the tool for work-related tasks.
LLM Setup. Participants accessed GPT-4 (GPT-4-0613) via a browser-based interface to an internally hosted “LLM-Gateway”, [^1] https://github.com/eSentire-Labs/LLM-Gateway, which provided a control and monitoring layer for LLM usage. The gateway enabled API-based access to OpenAI’s GPT-4 Enterprise platform under a 30-day data deletion policy and a non-training agreement in place. Analysts were free to navigate to the system and use the model to support their work tasks. The model was used without fine-tuning and had no access to the internet or internal SOC systems. Although other models were available, this study focuses exclusively on GPT-4.
Data Sources. Analysts used proprietary and third-party tools to access telemetry from endpoints, networks, identities, and application layers, including EDR, NDR, SIEM/XDR, firewalls, DNS, email, identity, cloud, WAF/CDN, and SaaS applications. Participants were permitted to use only Public and Internal data, with strict policies prohibiting submission of Confidential or Highly Confidential information [^2] Data was classified into four categories based on organisational sensitivity: Public data is freely accessible; Internal data comprises non-sensitive organisational information; Confidential data is protected by legal or contractual obligations; and Highly Confidential data is subject to the most stringent controls. Aside from this restriction, analysts were free to engage with the LLM as they saw fit. As the study focused on initial LLM engagement, all data were anonymised and devoid of personal identifiers. Accordingly, we report only aggregate usage trends, without attributing interactions to individual roles or experience levels.
@@ 4.2 Exploratory Analysis
Figure 3 (left) summarises analyst engagement with the LLM over 10 months. Usage was uneven, with a small number of analysts generating the majority of queries.
Given the skewed usage, we clustered the analysts based on their usage before analysing their long-term reliance on the LLM. We applied the Fisher-Jenks natural breaks algorithm [^3] https://en.wikipedia.org/wiki/Jenks_natural_breaks_optimization to total query counts per analyst. This generated three distinct clusters: a large group of low-usage analysts (Cluster 0, $n=34$), a smaller group of moderate users (Cluster 1, $n=10$), and a single power user (Cluster 2, $n=1$; most active analyst). Figure 3 (right) shows that daily usage increased steadily over time, driven by a subset of highly active analysts (a stronger increase was observed among 11 analysts), suggesting deeper integration of the LLM into daily workflows. The number of active analysts also grew month on month, suggesting sustained interest and continuous adoption of the LLM.
While a single power user submitted approximately 600 queries (17% of the total dataset), usage was not limited to a few individuals. By early 2024, all 45 analysts had interacted with the system, and daily query volumes increased from fewer than 10 queries per day in May-August 2023 to over 30 queries per day in January and February 2024. We analysed the monthly usage patterns and observed that the number of active analysts grew from 4 in May 2023 to 13 in June, reaching 18 by July and steadily increasing to a peak of 25 in February 2024. By early March 2024, 16 analysts had already submitted queries, showing strong ongoing engagement.
@@ 4.3 Query Level Analysis
This section discusses the thematic analysis results of individual queries. Each conversation could involve multi-step interactions. The thematic analysis was conducted on individual queries, rather than entire conversations, to retain fine-grained insights that a holistic approach might overlook. Section 4.4 presents conversation-level analysis. The appendix includes the list of all themes; full codebook can be accessed via the link in the footnote: [^4] https://osf.io/g8tr6/?view_only=db9c65552d364cba8c3a499bd12df589.
@@
The findings from the exploratory analysis reveal an interesting landscape of LLM use in our dataset. Despite wide variation in query volume, engagement style, and query lengths, usage trends point to a transition from periodic exploration to routine integration, with increased adoption over time and sustained interaction by a subset of analysts, with a stronger adoption among 11 analysts. These observations offer grounded insights into how LLMs are appropriated in practice, setting the stage for the next phase of our study: a qualitative analysis of the tasks and interaction strategies.
@@ 4.3.1 SOC Analyst Task Types
Task themes reflect the actual work that analysts aim to accomplish when interacting with an LLM. We find that analysts used LLMs for a range of tasks central to day-to-day SOC operations [1, 3]. These include interpreting and analysing commands; editing and improving the clarity of incident descriptions, documentation, and alert text; and understanding or troubleshooting code, scripts, and regular expressions. Analysts also used LLMs to analyse malware, explore system behaviours, build queries, and clarify email rules, highlighting the breadth of operational tasks. We highlight three common task themes below (see Figure 4 for the top 10), showing how LLMs integrate into workflows. For task adoption over time, see Figure 9 in the appendix.
@@
The LLM often began with an overview of the command, explaining the components and their relationships, inferring the intent behind complex operations, and highlighting suspicious behaviours. For example: “In summary, the [anonymised command] is trying to login as [anonymised user] without password and $\ldots$.
Many responses also included parameter breakdowns, example outcomes, risk implications, and recommended follow-up actions. Based on their sustained usage, we hypothesise that by decomposing command components and synthesising likely intent and behaviour, the LLM has shown the potential of enhancing the analysts’ situational understanding. This aligns with prior work (e.g., [17]), which finds that SOC analysts prefer explanations that contextualise evidence and inform action.
@@
In addition to the top three themes, analysts used LLMs for a wide range of other tasks including Tool/ Software/ System Understanding; Attack, Security & Threat Intelligence; Access Control & Authentication; File & Malware Analysis; Network & Traffic Analysis; and Query, Search & Data Processing. These tasks highlight specialised needs where LLMs offer targeted support. Notably, LLM Chat interactions reveal a growing rapport, with five analysts explicitly expressing thanks.